A rectangular corral has a perimeter of 64 yards. Its length is 8 yards more than its width. What are the dimensions of the corral?: A rectangular corral has a perimeter of 64 yards. Its length is 8 yards more than its width. What are the dimensions of the corral?
Answer by mangopeeler07(445)

x=width
x+8=length

2(x)+2(x+8)=64

Simplify
2x+2x+16=64

Combine like terms
4x+16=64

Subtract 16 from both sides
4x=48

Divide both sides by 4
x=12

width=12
length=20